An overhead wire stretches from the service drop that is 20ft from the ground,on a building to an insulator on a pole 35 ft from the ground. How long is the wire if the pole and building are 48 ft apart?(Assume ground is level.)Round answer to tenths.
Answer by nerdybill(7384)

An overhead wire stretches from the service drop that is 20ft from the ground,on a building to an insulator on a pole 35 ft from the ground. How long is the wire if the pole and building are 48 ft apart?(Assume ground is level.)Round answer to tenths.
.
The first thing you need to do is to draw a diagram!
.
From your diagram, you'll notice that the ground, building and pole forms a "right triangle". Because it is a right triangle (one of the interior angle of the triangle is 90 deg), we can apply Pythagorean theorem.
.
Let x = length of the overhead wire
then
x^2 = 48^2 + (35-20)^2
x^2 = 48^2 + 15^2
x^2 = 2304 + 225
x^2 = 2529
x = 50.3 ft